OpenStudy (anonymous):

What are the solutions of the quadratic equation? 4x^2 + 34x + 60 = 0 6 , 2/3 -5/2 , -2/3 -6 , -5/2 -6 , 4

OpenStudy (anonymous):

Well, the first step would be to simplify as much as possible before factoring. If you divide both sides by 2, you get: \[\large 2x^2 + 17x + 30 = 0\]

OpenStudy (anonymous):

Then, you try to factor. We know for certain it will look like this: \[\large (2x + ?)(x + ??) = 0\] Can you fill in the blanks? Remember: \[\large 2\cdot?? + 1\cdot ? = 17 \ \textrm{and} \\ \large ? \cdot ?? = 30\]

OpenStudy (anonymous):